# GraphKeeper: Graph Domain-Incremental Learning via Knowledge Disentanglement and Preservation

### v1

@@ 1 Introduction
Graph incremental learning (GIL)yuan2023continual ; febrinanto2023graph ; zhang2024continual aims to continuously update graph models as new graph data arrives, and has attracted increasing attention. Most existing methods target task-incremental (Task-IL) and class-incremental (Class-IL) settings (Figure 2), where new tasks or classes emerge within a single domain. However, as newly added graphs often come from different domains, the domain-incremental (Domain-IL) setting becomes essential. This is especially relevant with the rise of graph foundation models (GFMs), which require integrating diverse graphs from multiple domains to build a comprehensive and evolving knowledge base. Unfortunately, current GIL methods are designed for single-domain scenarios and struggle with Domain-IL. As illustrated in Figure 2, a representative method SSM zhang2022sparsified performs well under Class-IL but fails in the more challenging Domain-IL case. Despite its importance, effective GIL under Domain-IL remains an open problem.
@@
Refer to caption ((a)) Class-IL Scenario Refer to caption ((b)) Domain-IL Scenario Figure 2: Performance of SSM zhang2022sparsified , a representative GIL method, in Class-IL scenario and the more challenging Domain-IL scenario.
To tackle this problem, we first analyze the underlying cause of performance degradation in the Domain-IL setting, which is the catastrophic forgetting problem. In typical graph learning models such as GNNs, learning new graphs causes changes in model parameters that manifest in two aspects: (1) shifts in the embeddings of previously learned graphs, and (2) deviations in the decision boundary. These changes together lead to the forgetting of prior knowledge. Existing GIL methods attempt to constrain or adapt to such changes, which is relatively feasible in Task-IL and Class-IL scenarios, as the data still resides within a single domain. In contrast, Domain-IL involves substantial structural and semantic discrepancies across domains, making it much harder for GNNs to retain knowledge across diverse graph distributions. This difficulty is also evidenced by the negative transfer observed in multi-domain graph pre-training cao2023pre .
Based on the analysis above, we tackle the GIL challenges in Domain-IL scenario from two sides: (1) Embedding shifts: How to learn stable and disentangled representations for different graph domains in GIL? The embedding shifts introduced by model parameter changes bring the risk of semantic confusion between incremental graphs. To adapt to new domains, more drastic parameter changes occur, making embedding shifts more uncontrollable. (2) Decision boundary deviations: How to effectively retain knowledge from previously learned graph domains? The deviation in decision boundary is also a significant cause of catastrophic forgetting zheng2024learn , especially when the model adapts to new domains. Previous GIL methods, which treat embedding learning and prediction as an integrated process, are difficult to constrain the decision boundary effectively.
@@ 2.1 Graph Incremental Learning
Existing graph incremental learning methods can be categorized into regularization-based, memory replay-based, and parameter isolation-based. Regularization-based methods liu2021overcoming ; sun2023self ; cui2023lifelong aim to identify parameters that are important for previous tasks and constrain their changes through penalty terms. Memory replay-based methods zhou2021overcoming ; zhang2022sparsified ; wei2022incregnn ; wang2022streaming ; zhang2023ricci ; liu2023cat ; zhang2024topology ; niu2024graph are widely recognized for their effectiveness, explicitly preserving representative subsets of graphs from previous tasks and replaying them while learning new tasks. Parameter isolation-based methods zhang2022hierarchical ; zhang2023continual ; niu2024replay adapt to graph data in new tasks through adding trainable parameters. Despite the impressive performance in Task-IL and Class-IL scenarios, these methods do not consider the unique challenges in the Domain-IL scenario.
@@ 2.2 Multi-domain Graph Learning
Cross-domain Graph Pre-training. A promising approach for constructing GFMs is to pre-train on a large corpus of graphs, and transfer the pre-trained knowledge to downstream graphs zhang2023graph . Although existing cross-domain graph pre-training methods zhao2024all ; yu2024text ; xia2024anygraph ; zhao2024graphany can effectively adapt to downstream graphs, the model becomes fixed once pre-training is complete. Continuously learning knowledge from new graphs is essential zhang2024continual , especially when dealing with multiple downstream graph domains, but it faces the problem of catastrophic forgetting. Our method addresses the challenging Domain-IL scenario in GIL and can be integrated with various graph pre-training methods.
Graph Domain Adaptation. Graph domain adaptation (GDA) shi2025domain aims to leverage knowledge from the source domain to improve the performance in the target domain. Although multiple domains are considered, GDA and Domain-IL scenario in GIL are fundamentally different wu2024graph . For GDA, the source and target domains are simultaneously accessible in most cases, with the focus primarily on target domain performance. In comparison, for Domain-IL scenario, only the domain of current task is accessible, and the objective is to maintain performance across all previously learned domains.
@@ 4.1 Multi-domain Graph Disentanglement
where $\boldsymbol{F}_{i}$ is the original node features of $G_{i}$, $|G_{i}|$ denotes the number of nodes, $\operatorname{Proj}(\cdot)$ represents a projection operation, and $\overline{d}$ represents the unified feature projection dimension. The projection operation $\operatorname{Proj}(\cdot)$ is realized through truncated singular value decomposition (SVD) stewart1993early .
@@
Domain-specific Graph PEFT. To prevent potential catastrophic forgetting of previous domains due to drastic parameter changes when adapting to new domains, we aim to isolate parameters of different domains. Inspired by the success of low-rank adaptation (LoRA) hu2022lora ; li2024adaptergnn ; gui2024g ; yang2025graphlora , we propose graph domain-specific PEFT. Specifically, for graph domain in the task sequence, we equip the pre-trained GNN with a LoRA module. For the $i$-th domain, the $l$-th layer of the model $\psi_{i}$ is calculated as:
@@ 4.2 Deviation-Free Knowledge Preservation
Ridge Regression for Stable Classifier Updates. We aim to solve the decision boundary deviation issue by separating the classifier from the embedding model, thereby maintaining a stable decision boundary that remains unaffected by new incremental domains, which also aligns with the stable embeddings we introduce in the last section. To this end, we introduce a ridge regression-based knowledge preservation mechanism, which does not require gradient updates through back-propagation, thus avoiding the catastrophic forgetting caused by decision boundary deviations, inspired by zhuang2022acil ; zhuang2024gacl .
@@ 4.3 Domain-aware Distribution Discrimination
where $\hat{\boldsymbol{F}}$ denotes the transformed features. Though the $\mathrm{GNN}_{\text{proj}}$ is not trained, this high-dimensional random mapping sun2023feature ; mcdonnell2024ranpac helps to separate prototypes between different domains, particularly when utilizing the feature distance as the criterion for domain distribution discrimination. The confusion matrix of domain prototypes before and after mapping by Eq. (13) is shown in Figure E.2.
@@ 5.1 Experimental setup
Datasets and Baselines. To evaluate GraphKeeper [^1] https://github.com/RingBDStack/GraphKeeper, we conduct comprehensive experiments on 15 real-world datasets, where detailed descriptions are in Appendix C.1. We choose a variety of baselines, including general IL methods EWC kirkpatrick2017overcoming , MAS aljundi2018memory , GEM lopez2017gradient , LWF li2017learning and representative GIL methods TWP liu2021overcoming , ER-GNN zhou2021overcoming , SSM zhang2022sparsified , DeLoMe niu2024graph , PDGNNs zhang2024topology , TPP niu2024replay . Additionally, we include two self-designed baselines: Fine-Tune, which continuously training without any IL techniques, and Joint, which utilizes all previous graph domains for training.
Implementation Details. We implement all methods under the GIL benchmark zhang2022cglb . All results are averaged over 5 independent runs for a fair comparison. More details are provided in Appendix D.
@@ 5.2 Performance on Domain-IL Scenario
  • DeLoMe and PDGNNs exhibit relatively passable performance. These methods utilize SGC wu2019simplifying and APPNP gasteiger2018predict as their backbones, respectively, which sacrifices plasticity to mitigate forgetting. However, this limits their flexibility and applicability. We also try replacing their backbones as GCN and the results are summarized in Table E.3, showing a significant decline in performance.
@@ 5.3 Studies of Integration with GFMs
To further explore the potential of GraphKeeper, we integrate it with two representative GFMs, GCOPE zhao2024all and MDGPT yu2024text , to equip them with the continuous updating capability. Specifically, we first pre-train the GFMs and then incorporate GraphKeeper with them to evaluate the performance in the few-shot Domain-IL scenario. The results are summarized in Table 2.
@@ 5.5 Visualisation of Disentangled Embeddings
Embedding confusion is an important cause of catastrophic forgetting in Domain-IL scenario. To provide a more intuitive comparison of the advantages of GraphKeeper, we visualize the embeddings of different incremental graph domains with t-SNE van2008visualizing in Figure G.3 and Figure G.4.
@@ Acknowledgements
The corresponding author is Qingyun Sun. Authors of this paper are supported by the National Natural Science Foundation of China through grants No.62302023 and No.62225202, the Fundamental Research Funds for the Central Universities JK2024-07, and the Guangzhou Basic and Applied Basic Research Foundation 2024A04J3681. We owe sincere thanks to all authors for their valuable efforts and contributions.
@@ B.1 Algorithms
1
2Randomly initialize and frozen the parameters of $\mathrm{GNN}_{\text{proj}}$;
3Frozen the parameters of $\boldsymbol{\xi}$;
4Initialize the set $\boldsymbol{P}$ of the embedding prototypes;
5for $t=1,\dots,T$ do
@@
6
7  Align the feature dimensions of $G_{t}$ by Eq. (2);
8  Initialize the LoRA parameters $\boldsymbol{\phi}_{t}$ for $G_{t}$;
9 for $epoch=1,\dots,E$ do
10      Train the LoRA parameters $\boldsymbol{\phi}_{t}$ by Eq. (7);
11   end for
12
13  Calculate the embedding $\boldsymbol{X}_{t}$ by Eq. (3);
14  Add new embedding prototypes into set $\boldsymbol{P}$ by clustering;
@@
15
16  Update the ridge regression parameters $\boldsymbol{W}$ by Eq. (10) and (11);
17  Calculate the domain prototype $\boldsymbol{D}_{t}$;
18 end for
19
@@ 17  Calculate the domain prototype $\boldsymbol{D}_{t}$;
2 Calculate the test prototype $\boldsymbol{D}_{\text{test}}$;
3Discriminate the domain of $G_{\text{test}}$ by Eq. (14);
@@ 2 Calculate the test prototype $\boldsymbol{D}_{\text{test}}$;
4
5Calculate the embedding $\boldsymbol{X}_{\text{test}}$ with $\boldsymbol{\xi}$ and $\boldsymbol{\phi}_{\text{test}}$ by Eq. (3);
6Get the predicted result by Eq. (12);
@@ C.1 Datasets
We conduct experiments on 15 real-world datasets, including academic networks (Cora sen2008collective , Citeseer kipf2016semi , PubMed namata2012query , CoauthorCS shchur2018pitfalls , and DBLP fu2020magnn ), co-purchase networks (Photo shchur2018pitfalls and Computer shchur2018pitfalls ), web networks (WikiCS mernyei2020wiki , Facebook rozemberczki2021multi , Chameleon rozemberczki2021multi , and Squirrel rozemberczki2021multi ), social networks (GitHub rozemberczki2021multi , LastFMAsia rozemberczki2020characteristic , and DeezerEurope rozemberczki2020characteristic ), and airline networks (Airport chami2019hyperbolic ). Statistics of datasets are concluded in Table C.1. All the datasets are consented to by the authors for academic usage. All the datasets do not contain personally identifable information or offensive content.
@@ C.2.1 Domain-IL Scenario
Group 7
GitHub, LastFMAsia, DeezerEurope, WikiCS, Facebook,
Chameleon, Squirrel, Citeseer, Pubmed, CoauthorCS, DBLP
Group 8
Pubmed, Photo, WikiCS, Airport, CoauthorCS, Computer,
DeezerEurope, Cora, Facebook, LastFMAsia, Squirrel
@@ D.1 Implementation Details of GraphKeeper
We set the number of model layers to 2 for all methods, with the learning rate set to 5e-2, the weight decay coefficient set to 5e-4, and 200 training epochs per incremental graph domain. The parameters of models are optimized by Adam kingma2014adam . For GraphKeeper, the pre-trained GNN model is trained through link prediction on the first incremental domain (except for Sec. 5.3), the feature and structure augmentation are achieved by randomly masking a few features and dropping a few edges, and the embedding prototype sampling is implemented through DBSCAN ester1996density . The hyperparameter $r$ is chosen from {4, 8, 16, 32, 64, 128}, $\gamma_{1}$ is chosen from {0.01, 0.1, 1.0, 5.0, 10.0}, $\gamma_{2}$ is chosen from {0.001, 0.01, 0.10, 0.50, 1.0}.

### v2

@@ 1 Introduction
Graph incremental learning (GIL)Yuan et al. (2023); Febrinanto et al. (2023); Zhang et al. (2024b) aims to continuously update graph models as new graph data arrives, and has attracted increasing attention. Most existing methods target task-incremental (Task-IL) and class-incremental (Class-IL) settings (Figure 2), where new tasks or classes emerge within a single domain. However, as newly added graphs often come from different domains, the domain-incremental (Domain-IL) setting becomes essential. This is especially relevant with the rise of graph foundation models (GFMs), which require integrating diverse graphs from multiple domains to build a comprehensive and evolving knowledge base. Unfortunately, current GIL methods are designed for single-domain scenarios and struggle with Domain-IL. As illustrated in Figure 2, a representative method SSM Zhang et al. (2022c) performs well under Class-IL but fails in the more challenging Domain-IL case. Despite its importance, effective GIL under Domain-IL remains an open problem.
@@
Refer to caption ((a)) Class-IL Scenario Refer to caption ((b)) Domain-IL Scenario Figure 2: Performance of SSM Zhang et al. (2022c), a representative GIL method, in Class-IL scenario and the more challenging Domain-IL scenario.
To tackle this problem, we first analyze the underlying cause of performance degradation in the Domain-IL setting, which is the catastrophic forgetting problem. In typical graph learning models such as GNNs, learning new graphs causes changes in model parameters that manifest in two aspects: (1) shifts in the embeddings of previously learned graphs, and (2) deviations in the decision boundary. These changes together lead to the forgetting of prior knowledge. Existing GIL methods attempt to constrain or adapt to such changes, which is relatively feasible in Task-IL and Class-IL scenarios, as the data still resides within a single domain. In contrast, Domain-IL involves substantial structural and semantic discrepancies across domains, making it much harder for GNNs to retain knowledge across diverse graph distributions. This difficulty is also evidenced by the negative transfer observed in multi-domain graph pre-training Cao et al. (2023).
Based on the analysis above, we tackle the GIL challenges in Domain-IL scenario from two sides: (1) Embedding shifts: How to learn stable and disentangled representations for different graph domains in GIL? The embedding shifts introduced by model parameter changes bring the risk of semantic confusion between incremental graphs. To adapt to new domains, more drastic parameter changes occur, making embedding shifts more uncontrollable. (2) Decision boundary deviations: How to effectively retain knowledge from previously learned graph domains? The deviation in decision boundary is also a significant cause of catastrophic forgetting Zheng et al. (2024), especially when the model adapts to new domains. Previous GIL methods, which treat embedding learning and prediction as an integrated process, are difficult to constrain the decision boundary effectively.
@@ 2.1 Graph Incremental Learning
Existing graph incremental learning methods can be categorized into regularization-based, memory replay-based, and parameter isolation-based. Regularization-based methods Liu et al. (2021); Sun et al. (2023b); Cui et al. (2023) aim to identify parameters that are important for previous tasks and constrain their changes through penalty terms. Memory replay-based methods Zhou and Cao (2021); Zhang et al. (2022c); Wei et al. (2022); Wang et al. (2022); Zhang et al. (2023b); Liu et al. (2023); Zhang et al. (2024a); Niu et al. (2024b) are widely recognized for their effectiveness, explicitly preserving representative subsets of graphs from previous tasks and replaying them while learning new tasks. Parameter isolation-based methods Zhang et al. (2022b, 2023a); Niu et al. (2024a) adapt to graph data in new tasks through adding trainable parameters. Despite the impressive performance in Task-IL and Class-IL scenarios, these methods do not consider the unique challenges in the Domain-IL scenario.
@@ 2.2 Multi-domain Graph Learning
Cross-domain Graph Pre-training. A promising approach for constructing GFMs is to pre-train on a large corpus of graphs, and transfer the pre-trained knowledge to downstream graphs Zhang et al. (2023c). Although existing cross-domain graph pre-training methods Zhao et al. (2024a); Yu et al. (2024); Xia and Huang (2024); Zhao et al. (2024b) can effectively adapt to downstream graphs, the model becomes fixed once pre-training is complete. Continuously learning knowledge from new graphs is essential Zhang et al. (2024b), especially when dealing with multiple downstream graph domains, but it faces the problem of catastrophic forgetting. Our method addresses the challenging Domain-IL scenario in GIL and can be integrated with various graph pre-training methods.
Graph Domain Adaptation. Graph domain adaptation (GDA) Shi et al. (2025) aims to leverage knowledge from the source domain to improve the performance in the target domain. Although multiple domains are considered, GDA and Domain-IL scenario in GIL are fundamentally different Wu et al. (2024). For GDA, the source and target domains are simultaneously accessible in most cases, with the focus primarily on target domain performance. In comparison, for Domain-IL scenario, only the domain of current task is accessible, and the objective is to maintain performance across all previously learned domains.
@@ 4.1 Multi-domain Graph Disentanglement
where $\boldsymbol{F}_{i}$ is the original node features of $G_{i}$, $|G_{i}|$ denotes the number of nodes, $\operatorname{Proj}(\cdot)$ represents a projection operation, and $\overline{d}$ represents the unified feature projection dimension. The projection operation $\operatorname{Proj}(\cdot)$ is realized through truncated singular value decomposition (SVD) Stewart (1993).
@@
Domain-specific Graph PEFT. To prevent potential catastrophic forgetting of previous domains due to drastic parameter changes when adapting to new domains, we aim to isolate parameters of different domains. Inspired by the success of low-rank adaptation (LoRA) Hu et al. (2022); Li et al. (2024); Gui et al. (2024); Yang et al. (2025), we propose graph domain-specific PEFT. Specifically, for graph domain in the task sequence, we equip the pre-trained GNN with a LoRA module. For the $i$-th domain, the $l$-th layer of the model $\psi_{i}$ is calculated as:
@@ 4.2 Deviation-Free Knowledge Preservation
Ridge Regression for Stable Classifier Updates. We aim to solve the decision boundary deviation issue by separating the classifier from the embedding model, thereby maintaining a stable decision boundary that remains unaffected by new incremental domains, which also aligns with the stable embeddings we introduce in the last section. To this end, we introduce a ridge regression-based knowledge preservation mechanism, which does not require gradient updates through back-propagation, thus avoiding the catastrophic forgetting caused by decision boundary deviations, inspired by Zhuang et al. (2022, 2024).
@@ 4.3 Domain-aware Distribution Discrimination
where $\hat{\boldsymbol{F}}$ denotes the transformed features. Though the $\mathrm{GNN}_{\text{proj}}$ is not trained, this high-dimensional random mapping Sun et al. (2023a); McDonnell et al. (2024) helps to separate prototypes between different domains, particularly when utilizing the feature distance as the criterion for domain distribution discrimination. The confusion matrix of domain prototypes before and after mapping by Eq. (13) is shown in Figure E.2.
@@ 5.1 Experimental setup
Datasets and Baselines. To evaluate GraphKeeper [^1] https://github.com/RingBDStack/GraphKeeper, we conduct comprehensive experiments on 15 real-world datasets, where detailed descriptions are in Appendix C.1. We choose a variety of baselines, including general IL methods EWC Kirkpatrick et al. (2017), MAS Aljundi et al. (2018), GEM Lopez-Paz and Ranzato (2017), LWF Li and Hoiem (2017) and representative GIL methods TWP Liu et al. (2021), ER-GNN Zhou and Cao (2021), SSM Zhang et al. (2022c), DeLoMe Niu et al. (2024b), PDGNNs Zhang et al. (2024a), TPP Niu et al. (2024a). Additionally, we include two self-designed baselines: Fine-Tune, which continuously training without any IL techniques, and Joint, which utilizes all previous graph domains for training.
Implementation Details. We implement all methods under the GIL benchmark Zhang et al. (2022a). All results are averaged over 5 independent runs for a fair comparison. More details are provided in Appendix D.
@@ 5.2 Performance on Domain-IL Scenario
  • DeLoMe and PDGNNs exhibit relatively passable performance. These methods utilize SGC Wu et al. (2019) and APPNP Gasteiger et al. (2018) as their backbones, respectively, which sacrifices plasticity to mitigate forgetting. However, this limits their flexibility and applicability. We also try replacing their backbones as GCN and the results are summarized in Table E.3, showing a significant decline in performance.
@@ 5.3 Studies of Integration with GFMs
To further explore the potential of GraphKeeper, we integrate it with two representative GFMs, GCOPE Zhao et al. (2024a) and MDGPT Yu et al. (2024), to equip them with the continuous updating capability. Specifically, we first pre-train the GFMs and then incorporate GraphKeeper with them to evaluate the performance in the few-shot Domain-IL scenario. The results are summarized in Table 2.
@@ 5.5 Visualisation of Disentangled Embeddings
Embedding confusion is an important cause of catastrophic forgetting in Domain-IL scenario. To provide a more intuitive comparison of the advantages of GraphKeeper, we visualize the embeddings of different incremental graph domains with t-SNE Van der Maaten and Hinton (2008) in Figure G.3 and Figure G.4.
@@ Acknowledgements
The corresponding author is Qingyun Sun. Authors of this paper are supported by the National Natural Science Foundation of China through grants No.62302023 and No.62225202, the Fundamental Research Funds for the Central Universities JK2024-07, the Academic Excellence Foundation of BUAA for PhD Students, and the Guangzhou Basic and Applied Basic Research Foundation 2024A04J3681. We owe sincere thanks to all authors for their valuable efforts and contributions.
@@ B.1 Algorithms
Randomly initialize and frozen the parameters of $\mathrm{GNN}_{\text{proj}}$;
Frozen the parameters of $\boldsymbol{\xi}$;
Initialize the set $\boldsymbol{P}$ of the embedding prototypes;
for $t=1,\dots,T$ do
@@
  Align the feature dimensions of $G_{t}$ by Eq. (2);
  Initialize the LoRA parameters $\boldsymbol{\phi}_{t}$ for $G_{t}$;
 for $epoch=1,\dots,E$ do
      Train the LoRA parameters $\boldsymbol{\phi}_{t}$ by Eq. (7);
   end for
  Calculate the embedding $\boldsymbol{X}_{t}$ by Eq. (3);
  Add new embedding prototypes into set $\boldsymbol{P}$ by clustering;
@@
  Update the ridge regression parameters $\boldsymbol{W}$ by Eq. (10) and (11);
  Calculate the domain prototype $\boldsymbol{D}_{t}$;
end for
@@
Calculate the test prototype $\boldsymbol{D}_{\text{test}}$;
Discriminate the domain of $G_{\text{test}}$ by Eq. (14);
@@
Calculate the embedding $\boldsymbol{X}_{\text{test}}$ with $\boldsymbol{\xi}$ and $\boldsymbol{\phi}_{\text{test}}$ by Eq. (3);
Get the predicted result by Eq. (12);
@@ C.1 Datasets
We conduct experiments on 15 real-world datasets, including academic networks (Cora Sen et al. (2008), Citeseer Kipf and Welling (2017), PubMed Namata et al. (2012), CoauthorCS Shchur et al. (2018), and DBLP Fu et al. (2020)), co-purchase networks (Photo Shchur et al. (2018) and Computer Shchur et al. (2018)), web networks (WikiCS Mernyei and Cangea (2020), Facebook Rozemberczki et al. (2021), Chameleon Rozemberczki et al. (2021), and Squirrel Rozemberczki et al. (2021)), social networks (GitHub Rozemberczki et al. (2021), LastFMAsia Rozemberczki and Sarkar (2020), and DeezerEurope Rozemberczki and Sarkar (2020)), and airline networks (Airport Chami et al. (2019)). Statistics of datasets are concluded in Table C.1. All the datasets are consented to by the authors for academic usage. All the datasets do not contain personally identifable information or offensive content.
@@ C.2.1 Domain-IL Scenario
Group 7 GitHub, LastFMAsia, DeezerEurope, WikiCS, Facebook, Chameleon, Squirrel, Citeseer, Pubmed, CoauthorCS, DBLP
Group 8 Pubmed, Photo, WikiCS, Airport, CoauthorCS, Computer, DeezerEurope, Cora, Facebook, LastFMAsia, Squirrel
@@ D.1 Implementation Details of GraphKeeper
We set the number of model layers to 2 for all methods, with the learning rate set to 5e-2, the weight decay coefficient set to 5e-4, and 200 training epochs per incremental graph domain. The parameters of models are optimized by Adam Kingma and Ba (2014). For GraphKeeper, the pre-trained GNN model is trained through link prediction on the first incremental domain (except for Sec. 5.3), the feature and structure augmentation are achieved by randomly masking a few features and dropping a few edges, and the embedding prototype sampling is implemented through DBSCAN Ester et al. (1996). The hyperparameter $r$ is chosen from {4, 8, 16, 32, 64, 128}, $\gamma_{1}$ is chosen from {0.01, 0.1, 1.0, 5.0, 10.0}, $\gamma_{2}$ is chosen from {0.001, 0.01, 0.10, 0.50, 1.0}.